OpenStudy (anonymous):

Two triangles each have adjacent sides of length 120 ft and 180 ft. The first triangle has an angle between the two sides of 40°, while the second triangle has an angle between the two sides of 60°. What is the approximate difference between the areas of the two triangles?

OpenStudy (callisto):

Area = 1/2 (absinC) for the triangle with 40°, area = 1/2 (120x180xsin40°) = 6942.106185 sq.ft for the triangle with 60°, area = 1/2 (120x180xsin60°) = 9353.074361 sq.ft difference required = 9353.074361 - 6942.106185 = 2410.97 sq.ft

OpenStudy (anonymous):

thhank you callisto :)

OpenStudy (callisto):

welcome :)